News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 3 April 2015 9.00am EDT

During the last week the observing schedule was replanned to incorporate an observation of GK Persei, which was accepted as a Director's Discretionary Time Target of Opportunity on Mar 29. Loads for the replanned schedule will be uplinked later today. Scheduled observations of 2MASX J03534246+3714 and IRAS04124-0803 were impacted by the replan and will be rescheduled in future weeks.

An aspect camera dark current calibration was performed from the on-board loads on Mar 31 - Apr 1. An assessment of the data indicates nominal aspect camera performance and that the number of CCD warm pixels is increasing at the expected rate.

A real-time procedure was executed on Apr 1 to perform a routine self-check of the Electrical Interface Assembly (EIA) Sequencer.

The schedule of targets for the next two weeks is shown below and includes an observation of GRB 150314A, which is a Target of Opportunity follow-up accepted on Mar 19, an observation of RW Aur, which was accepted as a Director's Discretionary Time TOO on Mar 20, observations of the Crab and PG 1211+143 coordinated with HST and an observation of SDSS J1522+0327 coordinated with ground-based observatories.



------------------------------------------
Radiation Belts Apr 6 A1795 ACIS-S SwiftJ1644+57 ACIS-S A1795 ACIS-I Apr 7 BRC2 ACIS-I A1795 (7 obs) ACIS-I Radiation Belts Apr 8 PG1211+143 ACIS-S/HETG Apr 9 Crab ACIS-S Apr 10 PG1211+143 ACIS-S/HETG A1795 ACIS-S Apr 11 Radiation Belts SDSSJ1522+0327 ACIS-S Apr 12 PG1211+143 ACIS-S/HETG GRB150314A ACIS-S Apr 13 PG1211+143 ACIS-S/HETG Radiation Belts Apr 14 SgrA* ACIS-S SDSSJ080934.64+2548 ACIS-S A1795 ACIS-I Apr 15 PG1211+143 ACIS-S/HETG A1795 ACIS-S Apr 16 RWAur ACIS-S Radiation Belts PG1211+143 ACIS-S/HETG Apr 17 A1795 (3 obs) ACIS-I Apr 18 1928+738 ACIS-S Radiation Belts Apr 19

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports